When placed in water, sodium dichromate will dissociate in the following equation:
Na2Cr2O7--> 2Na + Cr2O7^2-

If the concentration of sodium dichromate is initially 0.127 mol/L, what will be the concentration of the sodium ion?
1
Expert's answer
2020-06-02T14:07:51-0400

[Na+]=C(Na2Cr2O7)*2=0.127*2=0.254 mol/L

Answer: 0.254 mol/L